Consider the Solow model with population growth and technological progress. The population grows at rate of d and the technology grows at rate of g. The depreciation rate of capital is λ. The aggregate production function is given as Y=100 ?![(1-u) ?]" where Y, K, L, ?, ? and u refers to aggregate output, aggregate capital stock, aggregate labor, output elasticity with respect to capital, output elasticity with

respect to labor, and natural rate of unemployment, respectively.

Draw a well-labeled graph that illustrates what happens to steady-state capital per worker, income per worker and consumption per worker over time in response to each of the following exogenous changes. (Hint: Plot per worker variables on y-axis against time on x-axis to show both immediate and gradual responses to the relevant exogenous change)

a.

The introduction of apprenticeship programs by the government

b.

A significant decline in the maintenance and repair expenses of machinery and equipment

c.

A permanent reduction in transportation costs
Answer #1

The solow growth model is an exogenous model of economic growth that analyzes changes in the level of output in an economy over time as a result of changes in the population growth rate, the saving rate, and the rate of technology progress.Assumptions of this model. 1.the population growth at a constant rate therefore current population represent by N and future population represent by N'are linked through the population growth equation N'=N(1+g).if the current population is 100 and the growth rate of population is 2 percent the future population is 102. 2.All consumers in the economy save a constant proportion 's'of their incomes and consume the rest. All firm in the economy produce output using same production technology that takes in capital and labour as inputs. Production function equation Y=F(K,L).The slow growth model assumes that the production function exhibits constant return to scale.solving the slow growth model in our analysis we assume that the production function takes the following from:Y=aKbL1-b where 0 greater than b and b greyer than 1.the income expenditure identity holds as an equilibrium condition:Y=c+i.consumer budget constraints Y=C+S. There fore in equilibrium i=S=sY.the solution concept used is that of a steady state the steady state is a state where the level of capital per worker does not change solving the following equation :K'=K=,>(1+g)k=(1-d)k+sakb.solow postulates a continuous production function linking output to the inputs of capital and labour which are substitubale slow models shows in his model that with variable technical coefficient there would be a tendency for capital labor ratio to adjust itself through time in the direction of equiblirium ratio. Slow analysis is convergent to equiblirium ratio. The steady state growth The concept of steady state growth is the counterpart of long run equilibrium in static theory. It is consistent with the concept of equilibrium growth. Taking different variables some of the new classical economists have given their interpretation to the concept of the steady state growth. Gw=Gn.joan Robinson describe the condition steady state growth as golden age of accumulation thus indicating a mythical state of affairs not likely to obtain in any actual economy.full employment in economy.two factors of production are capital and labour and they are paid according to their physical productivies. Labour and capital substitutable for each other. There is flexible system of price wage interest the model assume the production of a single composite commodity in the economy. Saving Investment for worker Balanced Growth investment Per worker, nk/N steady. State Point where saving is just enough to kSlow shows in his model that with variable technical coefficient there would be a tendency for capital labor ratio to adjust itself through time in the direction of equilibrium ratio. K=sY since output is produced with capital and labour, technology possibility are represented by the production function Y=F(K,L) that shows constant return to scale. In India the apprentice act was enacted 1961 agencies to recruit and develop well trained federal workers in highly skilled occupations.this site offers information andresources To Help federal agencies develop and maintain high performance apperenticship program. Packing materials movers drivers vehicles and possibly storage facilities.factors affect transport cost fuel cost and labour demand for freight.